The most common isotope of oxygen is oxygen-16 with a measured mass of 15.994915u.  

  1. Determine the mass defect;
  2. Determine the binding energy;
  3. Determine the binding energy per nucleon.
1
Expert's answer
2021-06-03T09:16:35-0400

The mass defect is "\Delta m = 8m_p + 8m_n + 8m_e - m_{\text{isotope}}" , because the atom consists of 8 protons, 8 neutrons and 8 electrons.

"\Delta m = 8(m_p + m_n + m_e) - m_{\text{isotope}}, \\\Delta m = 8(1.00728 + 1.00866+0.00055) - 15.994915, \\\Delta m = 0.137\,\mathrm{u} = 0.137\cdot 1.6605\cdot10^{-27} = 0.227\cdot10^{-27}\,\text{kg}."

The binding energy is "\Delta m c^2 = 0.227\cdot10^{-27}\cdot(3\cdot10^8)^2 = 2\cdot10^{-11}\,\mathrm{J}."

The binding energy per nucleon is "\dfrac{\Delta mc^2}{16} = 0.128\cdot10^{-11}\,\mathrm{J}."